How many grams of NH4NO3 are needed to produce 17.4 L of oxygen?
1
Expert's answer
2019-05-02T03:02:50-0400
"2NH_4NO_3(s)\rightarrow2N_2(g)+O_2(g)+4H_2O(g)""n=\frac{V}{V_m}"

"n(O_2)=\frac{V}{V_m}=\frac{17.4L}{22.4\frac{L}{mol}}=0.777 mol"

According to equation "n(NH_4NO_3):n(O_2)=2:1", then "n(NH_4NO_3)=n(O_2)\times2=0.777mol\times2=1.55 mol"

"m=M\times n"

"m(NH_4NO_3)=80.04\frac{g}{mol}\times 1.55mol= 124g"